# Jasmine: Harnessing Diffusion Prior for Self-Supervised Depth Estimation

### v1

@@ Abstract
In this paper, we propose Jasmine, the first Stable Diffusion (SD)-based self-supervised framework for monocular depth estimation, which effectively harnesses SD’s visual priors to enhance the sharpness and generalization of unsupervised prediction. Previous SD-based methods are all supervised since adapting diffusion models for dense prediction requires high-precision supervision. In contrast, self-supervised reprojection suffers from inherent challenges (e.g., occlusions, texture-less regions, illumination variance), and the predictions exhibit blurs and artifacts that severely compromise SD’s latent priors. To resolve this, we construct a novel surrogate task of hybrid image reconstruction. Without any additional supervision, it preserves the detail priors of SD models by reconstructing the images themselves while preventing depth estimation from degradation. Furthermore, to address the inherent misalignment between SD’s scale and shift invariant estimation and self-supervised scale-invariant depth estimation, we build the Scale-Shift GRU. It not only bridges this distribution gap but also isolates the fine-grained texture of SD output against the interference of reprojection loss. Extensive experiments demonstrate that Jasmine achieves SoTA performance on the KITTI benchmark and exhibits superior zero-shot generalization across multiple datasets.
@@ 1 Introduction
Estimating depth from monocular images is a fundamental problem in computer vision, which plays an essential role in various downstream applications such as 3D/4D reconstruction[48], autonomous driving[10], etc. Compared with supervised methods[32, 21], self-supervised monocular depth estimation (SSMDE) mines 3D information solely from video sequences, significantly reducing reliance on expensive ground-truth depth annotations. These methods derive supervision from geometric constraints (e.g., scene depth consistency) through cross-frame reprojection loss, and the ubiquitous video data further suggests an unlimited working potential. However, view reconstruction-based losses suffer from occlusions, texture-less regions, and illumination changes[56], which severely restrict the model’s capacity to recover fine-grained details and may cause pathological overfitting to specific datasets.
@@
Recent studies[24] demonstrated that SD possesses powerful visual priors to elevate depth prediction sharpness and generalization, which offers promising potential to address the above limitations. In addition, E2E FT[32] and Lotus[21] further reveal that single-step denoising can achieve better accuracy, which is particularly critical for self-supervised paradigms. It not only accelerates the inference denoising process but also significantly reduces the training costs in self-reprojection supervision, thereby creating opportunities to integrate SD into the SSMDE framework.
@@
However, fine-tuning diffusion models for dense prediction requires high-precision supervision to preserve their inherent priors[24]. Supervised methods typically employ synthetic RGB-D datasets, where clean depth annotations align with the high-quality SD’s training data, thereby keeping its latent space intact. In contrast, directly applying self-supervision introduces a critical challenge: reprojection losses or pre-trained depth pseudo-labels propagate perturbed gradients caused by artifacts and blurs into SD’s latent space, rapidly corrupting its priors during early training stages. Namely, high-precision “supervision” must exist at the beginning to protect SD’s latent space. Such supervision seems impossible in self-supervised learning, but we ingeniously find a handy and valuable alternative: the RGB image. In fact, the image inherently contains complete visual details, avoids external depth dependencies in self-supervision, and aligns perfectly with SD’s original objective of image generation. Therefore, we construct a surrogate task of hybrid image reconstruction through a task switcher, where the same SD model alternately reconstructs synthesized/real images and predicts depth maps within each training batch. This strategy repurposes self-supervised reprojection loss to tolerate color variations while maintaining structural consistency, intentionally decoupling color fidelity from depth accuracy, and finally preserving SD priors successfully.
@@
Another challenge is the output range of SD’s VAE that is inherently bounded within a fixed range, i.e., [-1, 1]. Existing methods typically normalize GT depth maps to this range in the training procedure. During inference, these supervised approaches perform least-squares alignment to recover absolute scale and shift, yielding scale- and shift-invariant (SSI) depth predictions. However, self-supervised frameworks rely on coupled depth-pose optimization, which theoretically requires shift invariance to be strictly zero for stable convergence, ultimately producing scale-invariant (SI) depth predictions. To bridge this inherent distribution gap, we propose a novel transform module termed Scale-Shift GRU (SSG). It not only iteratively aligns SSI depth to SI depth by refining scale-shift parameters but also acts as a gradient filter, which suppresses anomalous gradients caused by artifact-contaminated in self-supervised training, thereby preserving the fine-grained texture details of SD’s output while enforcing geometric consistency.
@@
Extensive experiments show that our proposed method, Jasmine, achieves the SoTA performance among all SSMDEs on the competitive KITTI dataset, shows remarkable zero-shot generalization across multiple datasets (even surpassing models trained with augmented data), and demonstrates unprecedented detail preservation. To sum up, the main contributions are summarized as follows:
@@
  • We proposed a surrogate task of HIR that anchors SD’s priors via self-supervised gradient sharing to avoid SD’s latent-space corruption caused by reprojection artifacts.
@@ 2 Relative Work
Due to the high costs of GT-depth collection from LiDARs and other sensors, self-supervised depth estimation (SSDE) has gained significant research attention. SSDE can be broadly categorized into stereo-based(learning depth from synchronized image pairs)[45, 14, 12, 3, 2] and monocular-based methods (using sequential video frames)[30, 60, 58, 31, 36, 56, 54, 20, 33, 19, 28]. Compared to the stereo ones, monocular approaches eliminate the need for synchronized binocular cameras and calibration, though they face challenges with dynamic objects and pose estimation inaccuracy. Additionally, when focusing on monocular video training, existing methods further diverge into single-frame and multi-frame approaches. Single-frame models neglect temporal information during inference, while multi-frame methods[46, 47, 17, 9, 1] leverage consecutive frames to construct cost volumes and even support iterative refinement at test time, yielding improved accuracy. However, the practical deployment of these methods remains constrained by the availability of multiple frames. Recently, the progress of DepthAnything v1/v2[50, 51] has revealed that, by training on large-scale image depth pairs, we can obtain an accurate single image depth prediction model with strong generalization. However, we argue that such datasets still remain a small fraction of the ubiquitous video data available. This observation motivates our exploration of the most challenging configuration: training exclusively on video sequences while maintaining single-frame inference capability, thereby laying the groundwork for developing genuinely versatile “DepthAnything” models.
@@
As the diffusion paradigm showcases its talents in generative tasks[22, 42, 27, 52], DDP[23] first reformulates depth perception as a depth map denoising task and leads to giant progress. Followers like DDVM[39], MonoDiffusion[41], and D4RD[44] (the latter two are self-supervised methods) all demonstrate the advantages of this paradigm in various MDE sub-tasks. Subsequently, the most renowned diffusion model, Stable Diffusion[38], has demonstrated significant potential for depth perception tasks. VPD[57], TAPD[25], and Prior-Diffusion[53] use Stable Diffusion as a multi-modal feature extractor, leveraging textual modality information to improve depth estimation accuracy. Concurrently, Marigold[24] and GeoWizard[11] enhanced model generalization and detail preservation by fine-tuning Stable Diffusion, capitalizing on its prior training with large-scale, high-quality datasets. Afterward, E2E FT[32] and Lotus[21] further accelerated inference by optimizing the noise scheduling process. In this work, our Jasmine continues these works and extends SD to the field of self-supervision.
Figure 2: The attempts to preserve the SD prior. The meanings of (a)-(f) are detailed in Sec. 3.2. Notably, while (e) demonstrates superior visual quality, it erroneously interprets surface textures (e.g., house windows) as depth edges. (g) shows the performance variations under different $\lambda$ settings for Ph supervision (Eq. 7) and latent supervision (Eq. 6). The complete metrics and their definitions are provided in the supplementary material.
@@ 3 Methods
In this section, we will introduce the foundational knowledge of the SSMDE and SD-based MDE (Sec. 3.1), the surrogate task of hybrid image reconstruction (Sec. 3.2), the scale-shift adaptation with GRU (Sec. 3.3) and the SD finetune protocol specified for self-supervision (Sec. 3.4). An overview of the whole framework is shown in Fig. 3.
Figure 3: Finetuning Protocol of Jasmine. The whole pipeline only contains 3 trainable modules: U-Net, Posenet, and Scale-Shift GRU (SSG). The other components are all frozen, including the encoder $\varepsilon$ and the decoder $\mathcal{D}$. In the beginning, the training image $\textbf{I}_{t}$ and the hybrid image $\textbf{I}_{h}$ are both concatenated with zero-noise $\mathbf{n}$[32] and fed into the $\varepsilon$. Next, the U-Net performs single-step denoising guided by the task switcher $s$, and subsequently decodes the depth prediction $D_{SSI}$ and the reconstructed image $\mathbf{I}$ with the $\mathcal{D}$ (Eq. 5). Afterward, the SSI-depth prediction is processed by the SSG for distribution refinement, yielding the final depth estimate $D_{SI}$. Supervision signals (right part) are detailed in Sec. 3.4.
@@ 3.1 Preliminaries
$L_{ph}=\eta_{p1}\frac{1-\operatorname{SSIM}\left(\textbf{I}_{t},\textbf{I}_{t^{\prime}\rightarrow t}\right)}{2}+\eta_{p2}\left\|\textbf{I}_{t}-\textbf{I}_{t^{\prime}\rightarrow t}\right\|.$ (2)
Stable Diffusion-based Monocular Depth Estimation reformulates depth prediction as an image-conditioned annotation generation task. Typically, given the image I and processed GT depth y, SD first encodes them to the low-dimension latent space through a VAE encoder $\varepsilon$, as $\textbf{z}^{\textbf{I}},\textbf{z}^{\textbf{y}}=\varepsilon(\textbf{I},\textbf{y})$. Afterward, gaussian noise is gradually added at levels $\tau\in[1,T]$ into $\textbf{z}^{\textbf{y}}$ to obtain the noisy sample, with $\mathbf{z^{y}_{\tau}}=\sqrt{\overline{\alpha}_{\tau}}\mathbf{z^{y}}+\sqrt{1-\overline{\alpha}_{\tau}}\mathbf{\epsilon},$ then the model learns to iteratively reverse it by removing the predicted noise:
and finally decodes the depth prediction with $D=\mathcal{D}(\mathbf{z_{0}^{y}})$. Here $\epsilon\sim\mathcal{N}(0,I),f^{\epsilon}_{\theta}$ is the $\epsilon$-prediction UNet, $\overline{\alpha}_{t}:=\prod_{s=1}^{t}(1-\beta_{s})$, and $\{\beta_{1},\beta_{2},\dots,\beta_{T}\}$ is the noise schedule with $T$ steps.
Equation 1 demonstrates that the self-supervised approach needs the depth prediction D for image warping. However, obtaining $\mathbf{z_{0}^{y}}$ requires iterative computation of Eq. 3, a process that becomes computationally infeasible considering the enormous size of SD models. Luckily, Lotus and E2E FT[32, 21] demonstrate that we can obtain comparable results with single-step denoising and directly predict depth,
$D=\mathcal{D}\left(f_{\theta}^{z}\left(\mathbf{z}^{\mathbf{y}}_{\tau},\mathbf{z^{I}}\right)\right),\tau=T,$ (4)
which makes it possible to train SD with self-supervision.
3.2 Surrogate Task: Image Reconstruction
Motivation.
The photometric reconstruction losses (Eq. 2) inevitably introduce the supervision with noise and artifacts due to occlusions, texture-less regions, and photometric inconsistencies. As shown in Fig. 2 (a), consider a scenario that $\mathcal{T}_{t\rightarrow t^{\prime}}$ represents a pure horizontal translation. The point $p$—which should ideally have 5 pixels of disparity (reciprocal of depth)—becomes occluded. It must instead displace 10 pixels to compensate for incorrect pixel matches, resulting in erroneous depth alignment with point $q$. This phenomenon propagates to neighboring points, collectively eroding structural details (e.g., the tree’s edge) while generating imprecise supervisory signals that rapidly degrade SD’s fine-grained prior knowledge.
To preserve these details, we notice that both SSMDE and SD inherently rely on image consistency: SSMDE uses photometric constraints (Eq. 2), while SD’s training directly minimizes image generation errors. Inspired by this, we propose a surrogate task: image reconstruction. Concretely, following [11], we also design a switcher $s\in\{s_{x},s_{y}\}$ to alternate the Unet $f_{\theta}^{z}$ between the main and surrogate tasks:
$[D,\mathbf{I}]=\mathcal{D}\left(f_{\theta}^{z}\left([s_{x},s_{y}],\mathbf{z}_{\tau}^{\mathbf{y}},\mathbf{z}^{\mathbf{I}}\right)\right).$ (5)
Therefore, we follow the SD paradigm and initially formulate the surrogate loss as:
$L_{s}=||\mathbf{z^{I}}-f_{\theta}^{z}(z^{\mathbf{I}},t,s_{y})||^{2}.$ (6)
Hybrid Images and Reconstruction Loss.
@@ 3.2 Surrogate Task: Image Reconstruction
We show that it is possible to preserve the SD priors by introducing a compact surrogate task. However, our experiments reveal that naively applying Eq. 6 for SSMDE optimization yields suboptimal results (Fig. 2(c)). Through empirical investigation, we identify three critical insights: 1) Reconstruction of inferior quality images (e.g., KITTI) introduces block artifacts (Fig. 2(c)). We attribute this to the latent space operating at $\frac{1}{8}$ resolution: When inputs align with the pre-trained $\varepsilon$ and $\mathcal{D}$, smooth supervision is achieved. Conversely, each latent pixel supervision manifests as 8×8 block artifacts in prediction. 2) Introducing high-quality synthesized images (maintaining self-supervision compliance) offers a potential solution, but exclusive reconstruction of synthetic data causes task decoupling and preserves depth prediction blurriness (Fig. 2 (d)). 3) Blending these images with a ratio $\lambda$ allows synthesized images to anchor the model to latent priors while real-world data enforces geometric structure alignment, is a possible solution. But this strategy shows notable sensitivity to $\lambda$ (Fig. 2 (g), (e) is a failure scene).
@@
To address this VAE-Image mismatch, we proposed to replace Eq. 6 with the photometric loss $L_{ph}$ (Eq. 2) in the image domain. Compared to Eq. 6, $L_{ph}$ emphasizes structural consistency rather than color fidelity, which aligns better with depth estimation objectives. As shown in Fig. 2(f, g), this supervision not only makes hybrid image reconstruction robust to $\lambda$ but also significantly improves the estimation quality. Therefore, we take the Hypersim dataset(4.2.1), a photorealistic synthetic dataset specifically designed for geometric learning, as the auxiliary image and update Eq. 6 with
$\mathbf{I_{h}}=R_{c}(\textbf{I}_{K},\textbf{I}_{H},\lambda),L_{s}=L_{ph}(\mathbf{I_{h}},\mathcal{D}(f_{\theta}^{z}(z^{\mathbf{I}},t,s_{y})),$ (7)
@@
where $R_{c}$ denotes random choice and $I_{K}$/$I_{H}$ is KITTI/Hypersim images.
@@
The specific synthetic data usage may raise concerns about applicability boundaries. To clarify, we conduct additional experiments and present three insights about the relationship between auxiliary data and performance: (1) Synthetic images are not essential, our surrogate task maintains efficacy with real-world imagery. (2) The dataset scale proves non-critical, as competitive performance emerges with samples under 1k. (3) Domain divergence between auxiliary and primary datasets enhances results and is even more important than image quality. Please refer to Sec. 4.3 for detailed experimental support.
This analysis reveals HIR is a highly promising training paradigm. It not only imposes no inherent limitations on any dense prediction tasks but also challenges the notion that fine-tuning SD requires high-quality annotation. Even with legacy datasets like KITTI, we can still leverage readily available images to effectively utilize SD priors and enhance depth estimation sharpness.

### v2

@@ Abstract
In this paper, we propose Jasmine, the first Stable Diffusion (SD)-based self-supervised framework for monocular depth estimation, which effectively harnesses SD’s visual priors to enhance the sharpness and generalization of unsupervised prediction. Previous SD-based methods are all supervised since adapting diffusion models for dense prediction requires high-precision supervision. In contrast, self-supervised reprojection suffers from inherent challenges (e.g., occlusions, texture-less regions, illumination variance), and the predictions exhibit blurs and artifacts that severely compromise SD’s latent priors. To resolve this, we construct a novel surrogate task of mix-batch image reconstruction. Without any additional supervision, it preserves the detail priors of SD models by reconstructing the images themselves while preventing depth estimation from degradation. Furthermore, to address the inherent misalignment between SD’s scale and shift invariant estimation and self-supervised scale-invariant depth estimation, we build the Scale-Shift GRU. It not only bridges this distribution gap but also isolates the fine-grained texture of SD output against the interference of reprojection loss. Extensive experiments demonstrate that Jasmine achieves SoTA performance on the KITTI benchmark and exhibits superior zero-shot generalization across multiple datasets. Project page and code are available at here.
@@ 1 Introduction
Estimating depth from monocular images is a fundamental problem in computer vision, which plays an essential role in various downstream applications such as 3D/4D reconstructiondepthsplat , autonomous drivingm4depth , etc. Compared with supervised methodsmarigold ; e2eft ; lotus , self-supervised monocular depth estimation (SSMDE) mines 3D information solely from video sequences, significantly reducing reliance on expensive ground-truth depth annotations. These methods derive supervision from geometric constraints (e.g., scene depth consistency) through cross-frame reprojection loss, and the ubiquitous video data further suggests an unlimited working potential. However, view reconstruction-based losses suffer from occlusions, texture-less regions, and illumination changesmonovit , which severely restrict the model’s capacity to recover fine-grained details and may cause pathological overfitting to specific datasets.
@@
Recent studiesmarigold demonstrated that SD possesses powerful visual priors to elevate depth prediction sharpness and generalization, which offers promising potential to address the above limitations. In addition, E2E FTe2eft and Lotuslotus further reveal that single-step denoising can achieve better accuracy, which is particularly critical for self-supervised paradigms. It not only accelerates the inference denoising process but also significantly reduces the training costs in self-reprojection supervision, thereby creating opportunities to integrate SD into the SSMDE framework.
@@
However, fine-tuning diffusion models for dense prediction requires high-precision supervision to preserve their inherent priorsmarigold . Supervised methods typically employ synthetic RGB-D datasets, where clean depth annotations align with the high-quality SD’s training data, thereby keeping its latent space intact. In contrast, directly applying self-supervision introduces a critical challenge: reprojection losses or pre-trained depth pseudo-labels propagate perturbed gradients caused by artifacts and blurs into SD’s latent space, rapidly corrupting its priors during early training stages. Namely, high-precision “supervision" must exist at the beginning to protect SD’s latent space. Such supervision seems impossible in self-supervised learning, but we find a handy and valuable alternative: the RGB image. In fact, the image inherently contains complete visual details, avoids external depth dependencies in self-supervision, and aligns perfectly with SD’s original objective of image generation. Therefore, we construct a surrogate task of mix-batch image reconstruction (MIR) through a task switcher, where the same SD model alternately reconstructs synthesized/real images and predicts depth maps within each training batch. This strategy repurposes self-supervised reprojection loss to tolerate color variations while maintaining structural consistencylrc , intentionally decoupling color fidelity from depth accuracy, and finally preserving SD priors successfully.
@@
Another challenge is the output range of SD’s VAEvae that is inherently bounded within a fixed range, i.e., [-1, 1]. Existing methods typically normalize GT depth maps to this range in the training procedure. During inference, these supervised approaches perform least-squares alignment to recover absolute scale and shift, yielding scale- and shift-invariant (SSI) depth predictions. However, self-supervised frameworks rely on coupled depth-pose optimization, which theoretically requires shift invariance to be strictly zero for stable convergence, ultimately producing scale-invariant (SI) depth predictions. To bridge this inherent distribution gap, we propose a gated recurrent unit (GRU)-based novel transform module termed Scale-Shift GRU (SSG). It not only iteratively aligns SSI depth to SI depth by refining scale-shift parameters but also acts as a gradient filter, which suppresses anomalous gradients caused by artifact-contaminated in self-supervised training, thereby preserving the fine-grained texture details of SD’s output while enforcing geometric consistency.
@@
Extensive experiments show that our proposed method, Jasmine, ❶ achieves the SoTA performance among all SSMDEs on the competitive KITTI dataset, ❷ shows remarkable zero-shot generalization across multiple datasets (even surpassing models trained with augmented data), ❸ demonstrates unprecedented detail preservation. As the first work to bridge self-supervised and zero-shot depth estimation paradigms, we also provide an in-depth analysis of the effects of different depth de-normalization strategies employed in their respective domains. To sum up, the main contributions are summarized as follows:
@@
  • We proposed a surrogate task of MIR that anchors SD’s priors via self-supervised gradient sharing to avoid SD’s latent-space corruption caused by reprojection artifacts.
@@ 2 Relative Work
Due to the high costs of GT-depth collection from LiDARs and other sensors, self-supervised depth estimation (SSDE) has gained significant research attention. SSDE can be broadly categorized into stereo-based(learning depth from synchronized image pairs)planedepth ; lrc ; ucnn ; pladenet ; falnet and monocular-based methods (using sequential video frames)hrdepth ; r-msfm ; devnet ; depthsegnet ; sd-ssmde ; monovit ; daccn ; gds ; rpr ; vifi . Additionally, when focusing on inference capability, existing methods can further diverge into single-frame and multi-frame approachesmanydepth ; depthformer ; dynamicdepth ; dualrefine . Comparisons between these paradigms are detailed in Sec. E.2. Recently, DepthAnything v1/v2dam1 ; dam2 has revealed that we can obtain an accurate single image depth prediction model with strong generalization by training on large-scale image depth pairs. However, we argue that such datasets still remain a small fraction of the ubiquitous video data available. This observation motivates our exploration of the most challenging configuration: training exclusively on video sequences while maintaining single-frame inference capability, thereby laying the groundwork for developing genuinely versatile “DepthAnything” models.
@@
As the diffusion paradigm showcases its talents in generative tasksddpm ; ddim ; liao2024denoising ; yue2024arbitrary , DDPDDP first reformulates depth perception as a depth map denoising task and leads to giant progress. Followers like DDVMddvm , MonoDiffusionmonodfs , and D4RDd4rd (the latter two are self-supervised methods but employ self-designed diffusion) all demonstrate the advantages of this paradigm in various MDE sub-tasks. Subsequently, the most renowned diffusion model, Stable Diffusionsd , has demonstrated significant potential for depth perception tasks. VPDvpd , TAPDtapd , and Prior-Diffusionpdfs use Stable Diffusion as a multi-modal feature extractor, leveraging textual modality information to improve depth estimation accuracy. Concurrently, Marigoldmarigold and GeoWizardgeo enhanced model generalization and detail preservation by fine-tuning Stable Diffusion, capitalizing on its prior training with large-scale, high-quality datasets. Afterward, E2E FTe2eft and Lotuslotus further accelerated inference by optimizing the noise scheduling process. In this work, our Jasmine continues these works and extends SD to the field of self-supervision.
@@ 3 Methods
In this section, we will introduce the foundational knowledge of the SSMDE and SD-based MDE (Sec. 3.1), the surrogate task of mix-batch image reconstruction (MIR, Sec. 3.2), the scale-shift adaptation with GRU (Sec. 3.3) and the SD finetune protocol specified for self-supervision (SSG, Sec. 3.4). An overview of the whole framework is shown in Fig. 2 and its training pseudocode is shown in Algorithm 1.
Figure 2: Finetuning Protocol of Jasmine. The $\textbf{I}_{t}$ and $\textbf{I}_{m}$ are each concatenated with $\mathbf{n}$e2eft and fed into the VAE encoder $\varepsilon$. Next, the U-Net performs single-step denoising guided by the task switcher $s$, and subsequently decodes the SSI-depth prediction $D_{SSI}$ and the reconstructed image with the $\mathcal{D}$ (Sec. 3.2). Afterward, the $D_{SSI}$ is processed by the SSG for distribution refinement, yielding the final depth estimation $D_{SI}$. The $L_{tc},L_{e},L_{ph}$ and $L_{s}$ are supervision loss and they are detailed in Sec. 3.4. The edge extraction module is detailed in Sec C
@@ 3.1 Preliminaries
where $\mathcal{T}_{t\rightarrow t^{\prime}}$ denotes the relative camera poses obtained from the pose network, $K$ denotes the camera intrinsics, and $\langle\cdot\rangle$ denotes the grid sample process. Then, we can compute the photometric reconstruction loss between $\textbf{I}_{t}$ and $\textbf{I}_{t^{\prime}\rightarrow t}$ to constrain the depth:
$L_{ph}(\textbf{I}_{t},\textbf{I}_{t^{\prime}\rightarrow t})=\eta_{p1}(1-\operatorname{SSIM}\left(\textbf{I}_{t},\textbf{I}_{t^{\prime}\rightarrow t}\right))/2+\eta_{p2}\left\|\textbf{I}_{t}-\textbf{I}_{t^{\prime}\rightarrow t}\right\|.$ (2)
Stable Diffusion-based Monocular Depth Estimation reformulates depth prediction as an image-conditioned annotation generation task. Typically, given the image I and processed GT depth y, SD first encodes them to the low-dimension latent space through a VAE encoder $\varepsilon$, as $(\textbf{z}^{\textbf{I}},\textbf{z}^{\textbf{y}})=\varepsilon(\textbf{I},\textbf{y})$. Afterward, gaussian noise is gradually added at levels $\tau\in[1,T]$ into $\textbf{z}^{\textbf{y}}$ to obtain the noisy sample, with $\mathbf{z^{y}_{\tau}}=\sqrt{\overline{\alpha}_{\tau}}\mathbf{z^{y}}+\sqrt{1-\overline{\alpha}_{\tau}}\mathbf{\epsilon},$ then the model learns to iteratively reverse it by removing the predicted noise:
and finally decodes the depth prediction with $D=\mathcal{D}(\mathbf{z_{0}^{y}})$. Here $\epsilon\sim\mathcal{N}(0,I),f^{\epsilon}_{\theta}$ is the $\epsilon$-prediction U-Net, $\overline{\alpha}_{t}:=\prod_{s=1}^{t}(1-\beta_{s})$, $\mathcal{D}$ is the VAE decoder and $\{\beta_{1},\beta_{2},\dots,\beta_{T}\}$ is the noise schedule with $T$ steps.
Equation 1 demonstrates that the self-supervised approach needs the depth prediction D for image warping. However, obtaining $\mathbf{z_{0}^{y}}$ requires iterative computation of Eq. 3, which becomes computationally infeasible considering the enormous size of SD models. Fortunately, Lotus and E2E FTe2eft ; lotus demonstrate that we can obtain comparable results with single-step denoising and directly predict depth, $D=\mathcal{D}\left(f_{\theta}^{z}\left(\mathbf{z}^{\mathbf{y}}_{\tau},\mathbf{z^{I}}\right)\right),\tau=T$, which makes it possible to train SD with self-supervision. The step-by-step workflow is detailed in Sec B.
3.2 Surrogate Task: Image Reconstruction
Self-supervision Compromise SD Prior.
The photometric reconstruction losses (Eq. 2) inevitably introduce the supervision with noise and artifacts due to occlusions, texture-less regions, and photometric inconsistencies. As shown in Fig. 3 (a), consider a scenario where relative camera poses $\mathcal{T}_{t\rightarrow t^{\prime}}$ represents a pure horizontal translation. The point pp—which should ideally have 5 pixels of disparity (reciprocal of depth)—becomes occluded. Instead, it must displace 10 to compensate for incorrect pixel matches, resulting in erroneous depth alignment with point $q$ (a detailed explanation is provided in Sec. F). This phenomenon propagates to neighboring points, collectively eroding structural details (e.g., the tree’s edge) while generating imprecise supervisory signals that rapidly degrade SD’s fine-grained prior knowledge.
To preserve these details, we notice that both SSMDE and SD inherently rely on image consistency: SSMDE uses photometric constraints (Eq. 2), while SD’s training directly minimizes image generation errors. Inspired by this, we propose a surrogate task: image reconstruction. Concretely, following geo , we design a switcher $s\in\{s_{x},s_{y}\}$ to alternate the U-Net $f_{\theta}^{z}$ between the main and surrogate tasks. When activated by $s_{x}$, we have $D=\mathcal{D}\left(f_{\theta}^{z}\left(s_{x},\mathbf{z}_{\tau}^{\mathbf{y}},\mathbf{z}^{\mathbf{I}}\right)\right)$. In contrast, we have $\mathbf{I}=\mathcal{D}\left(f_{\theta}^{z}\left(s_{y},\mathbf{z}_{\tau}^{\mathbf{y}},\mathbf{z}^{\mathbf{I}}\right)\right).$ Notably, the switcher $s$ is a processed one-hot vector and it is combined with the time embeddings fed into the $f_{\theta}^{z}$. This allows the U-Net to condition its internal operations on the currently selected task. Therefore, we follow the SD paradigm and initially formulate the surrogate loss as:
@@ 3.2 Surrogate Task: Image Reconstruction
Mix-batch Images Reconstruction with Photometric Supervision.
@@
We show that it is possible to preserve the SD priors by introducing a compact surrogate task. However, our experiments reveal that naively applying Eq. 4 for SSMDE optimization yields suboptimal results (Fig. 3(c)). Through empirical investigation, we identify three critical insights: 1) Inferior reconstructed images (e.g., KITTI) introduce block artifacts (Fig. 3(c)). We attribute this to the latent space operating at $\frac{1}{8}$ resolution: When inputs align with the pre-trained $\varepsilon$ and $\mathcal{D}$, smooth supervision is achieved. Conversely, each latent pixel supervision manifests as 8×8 block artifacts in prediction. 2) Introducing high-quality synthesized images (maintaining self-supervision compliance) offers a potential solution, but exclusive training on these data causes the model to only excel at reconstructing synthetic images but fails to generalize this capability to reduce depth estimation blurriness(Fig. 3 (d)). 3) Mixing these images within a training batch allows synthesized images to anchor the model to latent priors, while real-world data enforces geometric structure alignment, is a possible solution. But this strategy shows notable sensitivity to mix rate $\lambda$ (Fig. 3 (g), (e) is a failure scene).
@@
To address the mismatch between VAE and image, we proposed to replace Eq. 4 with the photometric loss $L_{ph}$ (Eq. 2) in the image domain. Compared to Eq. 4, $L_{ph}$ emphasizes structural consistency rather than color fidelity, which aligns better with depth estimation objectives. As shown in Fig. 3(f, g), this supervision not only makes MIR robust to $\lambda$ but also significantly improves the estimation quality. Therefore, we take the Hypersim dataset(4.2.1), a photorealistic synthetic dataset specifically designed for geometric learning, as the auxiliary image and update Eq. 4 with
$\mathbf{I_{h}}=\phi(\textbf{I}_{K},\textbf{I}_{H},\lambda),L_{s}=L_{ph}(\mathbf{I_{h}},\mathcal{D}(f_{\theta}^{z}(s_{y},\mathbf{z}_{\tau}^{\mathbf{y}},\mathbf{z^{I}}))),$ (5)
@@
where $\phi$ denotes random choice; $\textbf{I}_{K}$ and $\textbf{I}_{H}$ are KITTI and Hypersim images, respectively.
In summary, MIR constructs each training batch by randomly selecting images from these two datasets, and supervises the reconstructed image with the photometric loss in Eq. 5.
@@
The specific synthetic data usage may raise concerns about applicability boundaries. To clarify, we conduct additional experiments and present three insights about the relationship between auxiliary data and performance: (1) Synthetic images are not essential; our surrogate task maintains efficacy with real-world imagery. (2) The dataset scale proves non-critical, as competitive performance emerges with samples under 1k. (3) Domain divergence between auxiliary and primary datasets enhances results and is even more important than image quality. Please refer to Sec. 4.4 for detailed experimental support.
This analysis and related experiments reveal MIR is a highly promising training paradigm. It not only imposes no inherent limitations on any dense prediction tasks but also challenges the notion that fine-tuning SD requires high-quality annotation. Even with legacy datasets like KITTI, we can still leverage readily available images to effectively utilize SD priors and enhance depth estimation sharpness.
Figure 3: The attempts to preserve the SD prior. The meanings of (a)-(f) are detailed in Sec. 3.2. Notably, while (e) demonstrates superior visual quality, it erroneously interprets surface textures (e.g., house windows) as depth edges. (g) shows the performance variations under different $\lambda$ settings for photometric supervision (Eq. 5) and latent supervision (Eq. 4). The complete metrics and their definitions are provided in Sec. E.4.